# Chinese Academy of Sciences ${}^{3}$ Lab, Tencent PCG ${}^{4}$hanghai AI Laboratory https://github.com/Andrew0613/X-Restormer

### v3

@@ 1 Introduction
Image restoration aims to generate high-quality images from degraded images. In recent years, deep learning has achieved great success in this field, with numerous networks being proposed to address various image restoration tasks. Initially, networks are primarily designed to solve specific restoration tasks and are typically validated only on selected tasks. As deep learning techniques have continued to evolve, there has been an increasing focus on the development of general-purpose networks that can be applied to a broad range of tasks. This trend is particularly evident in the high-level vision field, where new backbone networks are being designed to support multiple tasks [17, 30], including classification, detection and segmentation. For image restoration, although more and more restoration backbones can handle multiple restoration tasks, their task generality is still limited, as illustrated in Footnote 2. For instance, SwinIR [25] achieves state-of-the-art performance on image super-resolution (SR) but falls short on image deblurring and dehazing. Conversely, Restormer [48] performs exceptionally well on image dehazing and deraining but is less effective on image SR. This discrepancy can be attributed to the fact that the characteristics of image degradation vary across different image restoration tasks. While all image restoration tasks involve mapping degraded images to clean images, the requirements for the capability of backbone networks differ depending on specific tasks.
@@ 2 Related Work
Image restoration networks. In the past years, numerous deep networks have been proposed for various image restoration tasks such as image SR [26, 59, 8], denoising [53, 54, 43], deblurring [1, 6], deraining [45, 4, 29] and dehazing [46, 41, 40]. Initially, most deep networks are designed for specific tasks. For instance, Zhang et al. [57] uses the channel attention to construct a CNN-based SR network. Zhang et al. [55] builds a denoising network by inputting multiple downsampled sub-images and a noise level map. Fu et al. [14] employs residual connections to design a deraining network. Cai et al. [3] estimates the transmission rate and inverts the atmospheric scattering model using a CNN to achieve image dehazing. Nah et al. [34] designs a multi-scale network for image deblurring.
@@
Recently, more and more methods have been developed to tackle multiple image restoration tasks. Zamir et al. [49] builds a multi-stage CNN for deraining, deblurring and denoising. Wang et al. [43] designs a U-shape Transformer for deraining, deblurring and denoising. Liang et al. [25] implements a Swin Transformer-based network that achieves state-of-the-art performance on SR, denoising and compression artifact reduction. Zamir et al. [48] proposes a novel transposed self-attention to build a U-shape network for deraining, deblurring and denoising. Chen et al. [5] constructs a U-shape CNN for denoising and deblurring. While existing methods have demonstrated some ability to generalize across several restoration tasks, their task generality remains limited.
@@ 4 General Backbone Network Design
Limitation of Restormer. In the benchmark experiments, Restormer shows inferior performance to SwinIR on SR, particularly on Urban100. The qualitative comparisons also indicate this phenomenon in Fig. 4. From the visual and LAM [16] results, We can observe that Restormer can exploit large-range and even global information for the reconstruction. However, compared to SwinIR, it fails to reconstruct fine textures, even for self-repeated patterns, This discrepancy can be attributed to the U-shape architecture adopted by Restormer on the one hand, which increases the difficulty of reconstructing high-frequency information. On the other hand, Restormer relies on depth-wise convolution for spatial information interaction, whose spatial mapping capability is relatively weaker than the spatial self-attention in SwinIR. Considering that the U-shape architecture is indispensable for some tasks, we still need to retain this architectural design for task generality. To overcome the limitation of Restormer and design a more powerful backbone network, we choose to further enhance its spatial information interaction ability. An intuitive and feasible solution is to incorporate the spatial self-attention module into Restormer.
@@ 5.2 Experimental Results
All-in-One Restoration. We conduct experiments on an all-in-one restoration setting to show the effectiveness of different backbone networks in addressing various tasks simultaneously. Networks are trained on five tasks with varying degradation levels (i.e., ×$\times 2$, ×$\times 4$ upscaling factors for SR and $\sigma\in(0,50)$ random noise level for Gaussian denoising). The sampling probability for each kind of task is the same, and the average performance of common benchmark datasets (appearing in Tab. 4 $\sim$ Tab. 7) is calculated. As shown in Tab. 8, with the relatively better task generality among the existing networks, Restormer exhibits better performance on the all-in-one restoration. By overcoming the limitation of Restormer, our X-Restormer demonstrates further advantages in handling multiple tasks concurrently, with its performance far exceeding other networks on all tasks. In contrast, the other networks are more or less affected by optimization conflicts across different tasks (e.g., SwinIR performs inferior to Restormer even on SR). These indicate that a general backbone network is of great significance for building a general model that process multiple image restoration tasks, which can effectively mitigate task conflicts with the performance drops.
Summary. With enhanced spatial mapping capability, our X-Restormer achieves significant performance improvement compared to Restormer. Specifically, X-Restormer obtains performance gains against Restormer of 0.42dB (Manga109), 0.22dB (Urban100), 0.54dB (HIDE), 0.61dB (Rain100H) and 0.93dB (SOTS Indoor) on image SR, denoising, deblurring, deraining and dehazing, respectively, showing the effectiveness of our design. Moreover, on all these five tasks, X-Restormer obtains state-of-the-art performance. Despite its simplicity, X-Restormer clearly has better task generality than other methods. Furthermore, we demonstrate that a powerful general backbone network has the potential to better handle multiple restoration tasks simultaneously. We hope it can inspire more works on the general image restoration backbone network design.

### v4

# Chinese Academy of Sciences 3Shanghai Artificial Intelligence Laboratory

@@ 1 Introduction
Image restoration aims to generate high-quality images from degraded images. In recent years, deep learning has achieved great success in this field, with numerous networks being proposed to address various image restoration tasks. Initially, networks are primarily designed to solve specific restoration tasks and are typically validated only on selected tasks. As deep learning techniques have continued to evolve, there has been an increasing focus on the development of general-purpose networks that can be applied to a broad range of tasks. This trend is particularly evident in the high-level vision field, where new backbone networks are being designed to support multiple tasks [17, 30], including classification, detection and segmentation. For image restoration, although more and more backbone networks can handle multiple restoration tasks, their task generality is still limited, as illustrated in Footnote 2. For instance, SwinIR [25] achieves state-of-the-art performance on image super-resolution (SR) but falls short on image deblurring and dehazing. Conversely, Restormer [48] performs exceptionally well on image dehazing and deraining but is less effective on image SR. This discrepancy can be attributed to the fact that the characteristics of image degradation vary across different image restoration tasks. While all image restoration tasks involve mapping degraded images to clean images, the requirements for the capability of backbone networks differ depending on specific tasks.
@@ 2 Related Work
Image restoration networks. In the past years, numerous deep networks have been proposed for various image restoration tasks such as image SR [26, 59, 8], denoising [53, 54, 43], deblurring [1, 6], deraining [45, 4, 29] and dehazing [46, 41, 40]. Initially, most deep networks are designed for specific tasks [57, 55, 14, 3, 34].
@@
Recently, with increasing attention to the task generality of networks, more and more methods have been developed to tackle multiple image restoration tasks. For instance, Zamir et al. [49] builds a multi-stage CNN for deraining, deblurring and denoising. Wang et al. [43] designs a U-shape Transformer for deraining, deblurring and denoising. Liang et al. [25] implements a Swin Transformer-based network that achieves state-of-the-art performance on SR, denoising and compression artifact reduction. Zamir et al. [48] proposes a novel transposed self-attention to build a U-shape network for deraining, deblurring and denoising. Chen et al. [5] constructs a U-shape CNN for denoising and deblurring. While existing methods have demonstrated some ability to generalize across several restoration tasks, their task generality remains limited.
@@ 4 General Backbone Network Design
Limitation of Restormer. In the benchmark experiments, Restormer shows inferior performance to SwinIR on SR, particularly on Urban100. The qualitative comparisons also indicate this phenomenon in Fig. 4. From the visual and LAM [16] results, We can observe that Restormer can exploit large-range and even global information for the reconstruction. However, compared to SwinIR, it fails to reconstruct fine textures, even for self-repeated patterns. This discrepancy can be attributed to the U-shape architecture adopted by Restormer on the one hand, which increases the difficulty of reconstructing high-frequency information. On the other hand, Restormer relies on depth-wise convolution for spatial information interaction, whose spatial mapping capability is relatively weaker than the spatial self-attention in SwinIR. Considering that the U-shape architecture is indispensable for some tasks, we still need to retain this architectural design for task generality. To overcome the limitation of Restormer and design a more powerful backbone network, we choose to further enhance its spatial information interaction ability. An intuitive and feasible solution is to incorporate the spatial self-attention module into Restormer.
@@ 5.2 Experimental Results
All-in-One Restoration. We conduct experiments on an all-in-one restoration setting to show the effectiveness of different backbone networks in addressing various tasks simultaneously. Networks are trained on five tasks with varying degradation levels (i.e., ×$\times 2$, ×$\times 4$ for SR and $\sigma\in(0,50)$ random level for denoising). The sampling probability for each task is the same, and the average performance on benchmark datasets is calculated. As shown in Tab. 8, with the relatively better task generality among the existing networks, Restormer exhibits better performance on the all-in-one restoration. By overcoming the limitation of Restormer, our X-Restormer demonstrates further advantages in handling multiple tasks concurrently, with its performance far exceeding other networks on all tasks. In contrast, the other networks are more or less affected by optimization conflicts across different tasks (e.g., SwinIR performs inferior to Restormer even on SR). These indicate that a general backbone network is of great significance for building a general model that process multiple image restoration tasks, which can effectively mitigate task conflicts with the performance drops.
Summary. With enhanced spatial mapping capability, our X-Restormer can significantly outperform Restormer. Specifically, X-Restormer obtains performance gains against Restormer of 0.42dB (Manga109), 0.22dB (Urban100), 0.54dB (HIDE), 0.61dB (Rain100H) and 0.93dB (SOTS Indoor) on image SR, denoising, deblurring, deraining and dehazing, respectively, showing the effectiveness of our design. Despite its simplicity, X-Restormer obtains state-of-the-art performance on all these five tasks and present the best task generality among the compared methods. Furthermore, we show that a more general backbone network can also better handle multiple restoration tasks simultaneously. We hope it can inspire more works on the general image restoration backbone network design.
@@ 6 Conclusion
Acknowledgements
This work was partially supported by National Natural Science Foundation of China (Grant No.62276251, 62272450), and the Joint Lab of CAS-HK. This work was also supported in part by Macau Science and Technology Development Fund under SKLIOTSC-2021-2023 and 0022/2022/A.